OpenStudy (anonymous):

A gas in a discharge tube consists entirely of ions of the same element, of the same charge. Each ion has only one remaining electron. The voltage between the electrodes of the gas discharge tube is increased from zero. Electrons are accelerated across the potential, and then hit the gaseous ions within the tube. These are called ballistic electrons. Below a threshold value of 91.9V the emission spectrum is completely blank. At a plate voltage of 91.9V the emission spectrum consists of a single line. Determine the chemical identity of the ion gas in the tube. Please enter the atomic number (Z

OpenStudy (anonymous):

E=q*v->1.6*10^-19*91.9=147.04*10^-19 then 147.04*10^-19=(1.6*10^-19)^2*8.987*10^9/r r=0.156*10^-10 ionic radius and i searched the periodic table but my answer wrong why????????

OpenStudy (anonymous):

3 i think